# Taming Data and Transformers for Audio Generation

### v1

@@ Abstract
Generating ambient sounds and effects is a challenging problem due to data scarcity and often insufficient caption quality, making it difficult to employ large-scale generative models for the task. In this work, we tackle the problem by introducing two new models. First, we propose AutoCap, a high-quality and efficient automatic audio captioning model. We show that by leveraging metadata available with the audio modality, we can substantially improve the quality of captions. AutoCap reaches CIDEr score of $83.2$, marking a $3.2\%$ improvement from the best available captioning model at four times faster inference speed. We then use AutoCap to caption clips from existing datasets, obtaining $761,000$ audio clips with high-quality captions, forming the largest available audio-text dataset. Second, we propose GenAu, a scalable transformer-based audio generation architecture that we scale up to 1.25B parameters and train with our new dataset. When compared to state-of-the-art audio generators, GenAu obtains significant improvements of $15.7\%$ in FAD score, $22.7\%$ in IS, and $13.5\%$ in CLAP score, indicating significantly improved quality of generated audio compared to previous works. This shows that the quality of data is often as important as its quantity. Besides, since AutoCap is fully automatic, new audio samples can be added to the training dataset, unlocking the training of even larger generative models for audio synthesis.
@@ 1 Introduction
Generative models have revolutionized the field of content creation, enabling the generation of high-quality natural images [91, 94, 92, 87, 36, 37], vivid videos [41, 111, 115, 89, 80], and intricate 3D shapes [10, 99, 60]. The domain of audio synthesis has undergone comparable advancement [43, 42, 65, 125, 32, 95, 83, 126, 25, 67, 116, 34], with three broad areas of study. Speech synthesis is the first, and probably the oldest problem in the domain, dating back to the eighteenth century [12, 118]. Modern text-to-speech synthesis adapts diffusion models used originally to synthesize images [50, 106, 33, 6]. The second domain is text-to-music synthesis, which has seen similar advances [2, 55, 68, 57, 88, 27, 84, 26, 134, 52, 133, 82]. The success in these two subdomains rests on two key pillars: (i) the availability of high-quality large-scale datasets containing text-to-speech, text-to-music data, and (ii) the development of scalable generative modeling methods [40, 101]. Indeed, there is an abundance of recordings of public speeches, performances, and musical pieces.
The world of audio synthesis is incomplete without the third category—ambient audio generation [65, 66, 43, 42, 29, 53], which is the main focus of this work. Unlike, the first two groups, ambient sound generation does not have a large corpus of accurately annotated samples readily available [46, 18]. It is fairly straightforward to collect text-to-speech data, by transcribing existing speeches. For ambient sounds, the task is substantially more challenging. Consider imagining a sound of “a peaceful early morning by the sea”, or the sound of “rocks falling on a wooden floor”. Going from an ambient sound to a caption is even more challenging as different sounds can suit the same visual scene. This is exploited in animation, where there are no real sounds to work with, and the entire sound scene is created artificially. The second challenge of ambient sound synthesis is the lack of available data. For visual modalities, such as images and videos [124, 81], huge datasets are collected from the Internet, filtered, processed, and annotated. Often such media is accompanied by raw description and metadata, which is used to train reliable automatic captioning models [8]. This is in contrast to the audio modality, where data sourced from video platforms like YouTube is often heavily edited or predominantly containing speech and music. Previous efforts to extract ambient sound clips from YouTube have shown a near $99\%$ rejection rate, making it impractical to compile a large-scale dataset.
In this work, we address the challenges of ambient sound synthesis from two sides. First, we introduce AutoCap, an efficient and high-quality audio captioning model that leverages visual information to reliably and automatically caption a large-scale dataset consisting of 761k audio clips. Second, we propose GenAu, a latent audio generator, based on a scalable FIT-architecture [80, 7].
@@
Current human-captioned audio datasets are limited, containing fewer than 51k text-audio pairs in total. This significantly impacts the training of current captioning models, making them more susceptible to overfitting and reducing their ability to generalize effectively. With this in mind, we build AutoCap, a state-of-the-art Automatic Audio Captioner (AAC). First, we refine the commonly used encoder-decoder transformer design based on a pretrained BART [58] model by introducing a Q-Former [59] that learns to summarize the encoded audio tokens into four times fewer tokens. By reducing the number of input tokens to the BART model, we speed up inference—an important step towards large-scale audio captioning—and provide better alignment with the original BART token representation due to the Q-Former additional capacity compared to simple projection layers used in previous work [49]. Second, we propose to use metadata and captions derived from video content to aid the captioning process and in this way, remedy the data scarcity problem affecting audio captioning methods. Critically, we augment the encoder inputs to assume both audio features and a set of descriptive textual metadata including audio title and a caption derived from the visual modality. This dual-input approach not only allows our model to achieve state-of-the-art performance on AudioCaps [46], marking a 3.2% improvement in CIDEr score, but it also helps reduce the domain gap with in-the-wild audios, improving model robustness. Equipped with this method, we obtain 761,000 audio clips from the major existing audio datasets and pair each audio sample with a descriptive synthetic caption. More importantly, since AutoCap is fully automatic, new audio samples can be easily added to the training set, paving the way to scaling the data beyond current constraints.
@@
To adapt audio generative models for larger scale training, we introduce GenAu, a scalable transformer-based architecture that achieves significant improvements over state-of-the-art audio generation models. Our approach introduces key architectural modifications over existing audio latent diffusion models [65, 43, 29, 42]. First, we train an efficient 1D-VAE [42] to transform a Mel-Spectrogram representation to a sequence of tokens and search for the optimal latent space for audio generation. Second, we recognize that audio grows fast temporally, and therefore, an efficient architecture that can handle larger sequences of tokens is needed. In particular, we employ a transformer architecture in the denoising backbone where we modify the FIT transformer [7] to generate audio in the latent space. Lastly, we extend the proposed FIT architecture to incorporate text conditioning through a dual encoder strategy. This involves an instruction-finetuned language model, FLAN-T5 [13], and an audio-centric CLAP encoding [49]. This adaptation significantly improves the model’s overall performance, achieving $15.7\%$ better FAD, $22.7\%$ higher Inception Score, and $13.5\%$ improvement in CLAP score, demonstrating superior audio-text alignment and audio generation quality.
@@
In summary, this work introduces: (i) AutoCap, a state-of-the-art audio captioner tailored towards the annotation of data at a large scale, which leverages audio metadata to improve accuracy and robustness, and a Q-Former to improve inference time and reduce overfitting; (ii) a large-scale dataset comprising 761k audio clips paired with synthetic captions derived from the proposed audio captioner and constituting the largest audio-text dataset currently available; (iii) GenAu, a novel audio generator based on a scalable transformer architecture specifically adapted to the audio domain. Our model achieves significantly improved quality when compared to the previous state-of-the-art.
@@ 2 Related Work
Automatic Audio Captioning (AAC). The goal of AAC is to produce natural language descriptions for given audio content. Most recent AAC methods [15, 119, 96, 102, 44, 14, 54, 123, 131] employ encoder-decoder transformer architectures, where an encoder receiving the audio signal produces a representation that is used by the decoder to produce the output caption. WavCaps [77] employs the CNN14 [51] and HTSAT [5] audio encoders and uses a pretrained BART [58] language decoder. CoNeTTE [136] proposes an audio encoder based on the ConvNeXt architecture and uses a vanilla transformer decoder [109] trained from scratch. Recently, EnCLAP [49] proposes the joint use of two audio representations in the form of CLAP [20] sequence embeddings and a discrete EnCodec [19] audio representation, and uses a pretrained BART model as the language backbone. Other work explores augmentation strategies to counter data scarcity [47, 136, 128]. Liu et al.[70] recently proposed to leverage the visual information using a pre-trained visual encoder to address sound ambiguities, reporting improvements. BART-Tags [31] generates captions conditioned on a sequence of predicted AudioSet tags. Our method uses audio metadata and visual information as additional conditioning signals and leverages a lightweight Q-Former [59] model that summarizes the audio feature to improve captioning speed and reduce model overfitting.
Text-Audio Datasets. The performance of text-audio models [135, 61, 16, 73, 17, 100, 21, 72, 103, 30, 11, 130], including AAC, is currently hindered by the lack of high-quality large-scale paired audio text data. The two main existing datasets are AudioCaps [46] and Clotho [18], comprising only 46k and 6k respectively of human-captioned audio clips. LAION-Audio [120] consists of 630k audio samples with raw descriptions, but annotation is highly noisy. WavCaps [77] proposes a filtering procedure based on ChatGPT [1] to collect 400k audio clips and weakly caption them based on the noisy descriptions alone. While weak-captioning does improve downstream metrics, it is suboptimal because it does not consider the audio signal. A recent work [43] explored a knowledge distillation approach that leverages data labels and a pre-trained audio captioner and retriever. In our work we collect 761k audio clips and produce captions using our state-of-the-art captioning method, resulting in a larger dataset with captions that are more aligned to the audio content.
@@
Text-conditioned audio generation. The current state-of-the-art text-to-audio generation methods widely adopt diffusion models [127, 53, 65, 66, 42, 29, 24, 112, 53]. AudioLDM [65] makes use of a latent diffusion model conditioned on CLAP embeddings, avoiding the need for the textual modality at training time. AudioLDM 2 [66] introduces a general representation of audio unifying the tasks of music, speech, and sound effects generation. Similarly, Audiobox [112] generates audio across different modalities such as speech and sound effects. StableAudio [24] introduces timing embeddings to allow the generation of long audios up to 95s. Recent work also explored controllable audio generation [97, 122, 79, 85, 132, 62, 64], visual-conditioned audio generation [117, 78, 113], and more recently join audio-video generation [104, 105, 121, 38, 107, 108, 3, 48, 114, 75, 9]. In this work, we show that improvements to data captioning quality and size, and the adoption of scalable architecture designs lead to state-of-the-art text-to-audio generation performance.
Figure 1: (Left) Overview of AutoCap. We employ frozen CLAP [49] and HTSAT [5] encoders to produce the audio representation. We then compact this representation into 4x fewer tokens using a Q-Former [59] module. This representation, along with tokens derived from pertinent metadata, is processed by a pretrained BART encoder-decoder model to generate the final caption.
@@
(Right) Overview of GenAu. Following latent diffusion models, we use a frozen 1D-VAE to convert a Mel-Spectrogram into latent sequences, which are then divided into groups and processed using ‘local’ attention layers based on the FIT architecture [7]. ‘Read’ and ‘write’ layers, implemented as cross-attention, facilitate information transfer between input latents and learnable latent tokens. Finally, ‘global’ attention layers on latent tokens allow for global communication across all groups.
@@ 3 Method
In this section, we describe our approach to high-quality text-to-audio generation. Sec. 3.1 describes our method for producing accurate audio captions using our AutoCap model, Sec. 3.2 describes our strategy for building our large-scale captioned audio dataset, including data selection and re-captioning, lastly Sec. 3.3 describes our scalable audio generation architecture for our GenAu model.
@@ 3.1 Automatic Audio Captioning
The availability of textual descriptions that are highly aligned with the modality to be generated, coupled with large-scale datasets is of critical importance in achieving high-quality generation results [98]. In this context, automatic image and video captioning have been successfully employed to improve text-to-image and text-to-video models by augmenting and improving the quality of the training data for these models [98, 8]. Despite the large quantity of available non-captioned audio data, current state-of-the-art Automatic Audio Captioning (AAC) methods have not been used to improve the quality of the training data of text-to-audio models [77, 136, 49]. We posit that this is in part due to the lower quality of the produced captions for existing models compared to similar models available for other modalities. AAC methods have generally adopted an encoder-decoder transformer design, where an audio encoder (e.g. [19, 51, 5]) is responsible for producing an output representation that can be taken as input by a decoder model to produce an output audio caption. While audio-text pairs are difficult to obtain from web data, audio files from many sources are still commonly associated with metadata that might be relevant for captioning such as raw user descriptions, or captions derived from a related modality (i.e. accompanied visual information). Motivated by this observation and the benefits of better-aligned captions, we propose AutoCap, an audio captioning model where the encoded audio representation is augmented with metadata. Fig. 1 presents an overview of AutoCap.
We consider a dataset of audio signals paired with a corresponding caption $\langle\mathbf{a},\mathbf{y}\rangle$ and metadata represented as a set of token sequences $\{\mathbf{m}_{j}\}_{j=1}^{j=M}$. Inspired by state-of-the-art AAC methods [77, 136, 49], we employ an encoder-decoder sequence-to-sequence model. We start by computing a global feature representation of the audio:
@@
where $\mathcal{P}_{\mathrm{clap}}$ is a learnable projection layer and $\mathcal{E}_{\mathrm{clap}}$ is the audio encoder of a pretrained CLAP model [20]. Then we compute a separate local feature representation of the input audio:
@@
where $\mathcal{Q}$ is a Q-Former [59] that outputs a compact sequence audio representation and $\mathcal{E}_{\mathrm{a}}$ is the HTSAT [5] audio encoder that produces a time-aligned representation. The Q-Former efficiently learns 256 latent tokens, which serve as keys in cross-attention layers with the input features, thereby condensing the audio input features into 256 tokens. Metadata sequences $\mathbf{m}_{i}$ are then embedded using the embedding layer of the pretrained encoder-decoder model to obtain corresponding embedding sequences $\mathbf{x}_{\mathrm{meta}_{i}}$. We represent the input audio and metadata as the following input sequence:
$\mathbf{x}=\mathbf{x}_{\mathrm{clap}}~{}\texttt{\small[boa]}~{}\mathbf{x}_{\mathrm{audio}}~{}\texttt{\small[eoa]}~{}\texttt{\small[bom]}_{1}~{}\mathbf{x}_{\mathrm{meta}_{1}}~{}\texttt{\small[bom]}_{1}~{}...~{}\texttt{\small[bom]}_{M}~{}\mathbf{x}_{\mathrm{meta}_{M}}~{}\texttt{\small[bom]}_{M}\text{,}$ (3)
@@
$\mathbf{\hat{y}}=\mathcal{D}_{\mathrm{t}}(\mathcal{E}_{\mathrm{t}}(\mathbf{x}))\text{,}$ (4)
where $\mathcal{E}_{\mathrm{t}}$ is an audio encoder that produces a representation of the input signal $\mathbf{x}$ which is fed to the decoder $\mathcal{D}_{\mathrm{t}}$. We adopt a pretrained BART transformer model [58] as our encoder-decoder. Finally, we train our model using a standard cross-entropy loss over next token predictions:

### v2

@@ Abstract
Generating ambient sounds is a challenging task due to data scarcity and often insufficient caption quality, making it difficult to employ large-scale generative models for the task. In this work, we tackle this problem by introducing two new models. First, we propose AutoCap, a high-quality and efficient automatic audio captioning model. By using a compact audio representation and leveraging audio metadata, AutoCap substantially enhances caption quality, reaching a CIDEr score of $83.2$, marking a $3.2\%$ improvement from the best available captioning model at four times faster inference speed. Second, we propose GenAu, a scalable transformer-based audio generation architecture that we scale up to 1.25B parameters. Using AutoCap to generate caption clips from existing audio datasets, we demonstrate the benefits of data scaling with synthetic captions as well as model size scaling. When compared to state-of-the-art audio generators trained at similar size and data scale, GenAu obtains significant improvements of $4.7\%$ in FAD score, $22.7\%$ in IS, and $13.5\%$ in CLAP score, indicating significantly improved quality of generated audio compared to previous works. Moreover, we propose an efficient and scalable pipeline for collecting audio datasets, enabling us to compile 57M ambient audio clips, forming AutoReCap-XL, the largest available audio-text dataset, at 90 times the scale of existing ones. Our code, model checkpoints, and dataset are publicly available.
@@ 1 Introduction
Text-conditioned generative models have revolutionized the field of content creation, enabling the generation of high-quality natural images (Ramesh et al., 2022; Rombach et al., 2022; Podell et al., 2023; Haji-Ali et al., 2024), vivid videos  Ho et al. (2022); Villegas et al. (2022); Wang et al. (2023b); Qiu et al. (2023); Menapace et al. (2024), and intricate 3D shapes (Cheng et al., 2023). The domain of audio synthesis has undergone comparable advancement (Huang et al., 2023b; a; Liu et al., 2023b; Xue et al., 2024; Guan et al., 2024; Saito et al., 2024; Niu et al., 2024; Yang et al., 2023a; Evans et al., 2024b; Liu et al., 2024; Wang et al., 2024b; Guo et al., 2023), with three broad areas of study: speech, music and ambient sounds. The success in these domains rests on two key pillars: (i) the availability of high-quality large-scale datasets with text annotations, and (ii) the development of scalable generative models (Ho et al., 2020; Song et al., 2020).
In the field of audio synthesis, ambient audio generation emerges as a critical domain, which is the main focus of this work. Unlike speech and music, ambient sound generation suffers from a lack of extensive, well-annotated datasets (Kim et al., 2019; Drossos et al., 2020). Attempts to curate ambient audio from online videos predominantly failed, primarily due to the dominance of speech and music content in such videos. For instance, AudioSet (Gemmeke et al., 2017), the largest available audio dataset sourced from online videos, contains $99\%$ speech or music clips. Previous efforts to filter ambient audio from similar datasets involved using expensive classifiers on the video or audio content, making it impractical to compile a large-scale dataset due to the high rejection rate. In this work, we propose a simple, yet scalable filtering approach that leverages existing automatic video transcription to identify segments with ambient sounds. This method is not only more efficient but also more feasible, as it eliminates the need to download the audio or video content. Additionally, by using time-aligned transcripts, we reduce the rejection rate to only $83\%$. Through this approach, we built AutoReCap-XL, a dataset containing 57 million ambient audio clips sourced from existing video datasets, representing a 90-fold increase over the size of previously available datasets.
Another challenge in compiling large-scale text-audio datasets is providing accurate textual descriptions. For visual modalities, such as images and videos (Xue et al., 2022; Miech et al., 2019), researchers often relied on the raw description and metadata to train strong visual-text models including reliable captioners (Chen et al., 2024b). Similarly, speech modality benefits from strong automatic transcription models to provide textual annotations. For ambient sounds, however, the task is substantially more challenging as accompanying raw text tends to describe visual information or convey feelings, rather than detailing the audio content. Moreover, human-captioned audio datasets are limited, containing fewer than $51k$ text-audio pairs in total. This significantly impacts the training of current captioning models, making them more susceptible to overfitting and reducing their ability to generalize effectively. In this work, we address this challenge by introducing AutoCap, an efficient and high-quality audio captioning model that leverages visual information to enhance captioning.
@@
AutoCap refines the commonly used encoder-decoder design based on a pretrained BART (Lewis et al., 2020) model by introducing a Q-Former (Li et al., 2023a) that learns to summarize the encoded audio tokens into four times fewer tokens. By reducing the number of input tokens to the BART model, we speed up inference—an important step towards large-scale audio captioning—and provide better alignment with the original BART token representation due to the Q-Former additional capacity compared to simple projection layers used in previous work (Kim et al., 2024b). Second, we propose to use metadata and captions derived from video content to aid the captioning process and in this way, remedy the data scarcity problem. Critically, we augment the encoder inputs to assume both audio features and a set of descriptive textual metadata including audio title and a caption derived from the visual modality. This dual-input approach not only allows our model to achieve state-of-the-art performance on AudioCaps (Kim et al., 2019), marking a 3.2% improvement in CIDEr score, but it also helps reduce the domain gap with in-the-wild audios.
@@
Moreover, to adapt audio generative models for larger scale training, we introduce GenAu, a scalable transformer-based architecture that achieves significant improvements over state-of-the-art audio generation models. Our approach introduces key architectural modifications over existing audio latent diffusion models (Liu et al., 2023b; Huang et al., 2023b; Ghosal et al., 2023; Huang et al., 2023a). First, we train an efficient 1D-VAE (Huang et al., 2023a) to transform a Mel-Spectrogram representation to a sequence of tokens and search for the optimal latent space for audio generation. Second, we recognize that audio grows fast temporally, yet contains many silent and redundant segments. Therefore, an efficient architecture that can handle such properties is needed. In particular, we employ a transformer architecture in the denoising backbone where we modify the FIT transformer (Chen & Li, 2023) to generate audio in the latent space. Lastly, we extend the proposed FIT architecture to incorporate text conditioning through a dual encoder strategy. This involves an instruction-finetuned language model, FLAN-T5 (Chung et al., 2022), and an audio-centric CLAP encoding (Kim et al., 2024b). This adaptations significantly improves the model’s performance over exiting methods, achieving $22.7\%$ higher Inception Score, $4.7\%$ better FAD, and $13.5\%$ improvement in CLAP score, demonstrating superior audio-text alignment and audio generation quality.
Finally, we explore the scaling behavior of text-to-audio diffusion models in relation to model size and data size. Prior studies in text-to-image generation have established a scaling trend where performance improves with the increases in data volume and model size (Peebles & Xie, 2023; Li et al., 2024). Such exploration, however, has not been sufficiently conducted for the audio modality. Initially, we analyze the impact of augmenting the dataset with synthetic captions on model performance. Our findings reveal a clear trend of improvement in FD and IS as we increase the amount of training data. Furthermore, we observe a consistent trend of enhanced performance across all metrics when scaling up the model size, concluding that the audio modality also benefits significantly from increases in both model size and data scale.
@@
In summary, this work introduces: (i) AutoCap, a state-of-the-art audio captioner tailored towards the annotation of data at a large scale, which leverages audio metadata to improve accuracy and robustness, and a Q-Former to improve inference time and reduce overfitting; (ii) GenAu, a novel audio generator based on a scalable transformer architecture specifically adapted to the audio domain. Our model achieves significantly improved quality when compared to the previous state-of-the-art. (iii) AutoReCap-XL, the largest available audio dataset, comprising 57M audio clips paired with synthetic captions derived from the proposed audio captioner.
@@ 2 Related Work
Automatic Audio Captioning (AAC). The goal of AAC is to produce natural language descriptions for given audio content. Most recent AAC methods (Deshmukh et al., 2023; Wu et al., 2024; Salewski et al., 2023; Sridhar et al., 2023; Kadlčík et al., 2023; Cousin et al., 2023; Labbé et al., 2023; Xu et al., 2023; Zhang et al., 2024) employ encoder-decoder transformer architectures, where an encoder receiving the audio signal produces a representation that is used by the decoder to produce the output caption. WavCaps (Mei et al., 2023a) employs the CNN14 (Kong et al., 2019) and HTSAT (Chen et al., 2022) audio encoders and uses a pretrained BART (Lewis et al., 2020) language decoder. CoNeTTE (Étienne Labbé et al., 2023) proposes an audio encoder based on the ConvNeXt architecture and uses a vanilla transformer decoder (Vaswani et al., 2017) trained from scratch. Recently, EnCLAP (Kim et al., 2024b) proposes the joint use of two audio representations in the form of CLAP (Elizalde et al., 2023) sequence embeddings and a discrete EnCodec (Défossez et al., 2022) audio representation and uses a pretrained BART model as the language backbone. Other work explores augmentation strategies to counter data scarcity (Kim et al., 2022; Étienne Labbé et al., 2023; Ye et al., 2022). Liu et al. (2023d) recently proposed to leverage the visual information using a pre-trained visual encoder to address sound ambiguities, reporting improvements. BART-Tags (Gontier et al., 2021) generates captions conditioned on a sequence of predicted AudioSet tags. Our method uses audio metadata and visual information as additional conditioning signals and leverages a lightweight Q-Former (Li et al., 2023a) model that summarizes the audio feature to improve captioning speed and reduce overfitting.
@@
Text-conditioned audio generation. The current state-of-the-art text-to-audio generation methods widely adopt diffusion models (Yang et al., 2023b; Kreuk et al., 2023; Liu et al., 2023b; c; Huang et al., 2023a; Ghosal et al., 2023; Evans et al., 2024a; Vyas et al., 2023; Kreuk et al., 2023). AudioLDM (Liu et al., 2023b) makes use of a latent diffusion model conditioned on CLAP embeddings, reducing the need for the textual modality at training time. AudioLDM 2 (Liu et al., 2023c) introduces a general representation of audio unifying the tasks of music, speech, and sound effects generation. Similarly, Audiobox (Vyas et al., 2023) generates audio across different modalities such as speech and sound effects. StableAudio (Evans et al., 2024a) introduces timing embeddings to allow the generation of long audios up to 95s. Recent work also explored controllable audio generation (Shi et al., 2023; Xu et al., 2024; Melechovsky et al., 2024; Paissan et al., 2024; Zhang et al., 2023b; Liang et al., 2024; Liu et al., 2023a), visual-conditioned audio generation (Wang et al., 2024c; Mei et al., 2023b; Wang et al., 2023a), and more recently joint audio-video generation (Tang et al., 2023a; b; Xing et al., 2024; Hayakawa et al., 2024; Tian et al., 2024; Vahdati et al., 2024; Chen et al., 2024a; Kim et al., 2024a; Wang et al., 2024a; Mao et al., 2024; Chen et al., 2024c). In this work, we show that improvements to data captioning quality and size, and the adoption of scalable architecture designs lead to state-of-the-art text-to-audio generation performance.
Text-Audio Datasets. The performance of text-audio models (Zhu et al., 2024; Li et al., 2023b; Deshmukh et al., 2024a; Mahfuz et al., 2023; Deshmukh et al., 2024b; Shu et al., 2023; Elizalde et al., 2024; Liu et al., 2023f; Tang et al., 2024; Gong et al., 2024; Cheng et al., 2024; Zhang et al., 2023a), including AAC, is currently hindered by the lack of high-quality large-scale paired audio text data of ambient sounds. The two main existing datasets are AudioCaps (Kim et al., 2019) and Clotho (Drossos et al., 2020), comprising only $46k$ and $6k$ respectively of human-captioned audio clips. Another challenge is the limited availability of audio clips from sound-only platforms. LAION-Audio (Wu et al., 2023) relied on numerous sources of audio platforms such as BBC Sound Effects (BBC Sound Effects, 2024),  (Font et al., 2013) FreeSounds, and SoundBible (SoundBible, 2024) to form a dataset consisting of 630k audio samples with highly noisy raw descriptions. WavCaps (Mei et al., 2023a) proposes a filtering procedure based on ChatGPT (Achiam et al., 2023) to collect $400k$ audio clips and weakly caption them based on the noisy descriptions alone. Several subsequent work (Majumder et al., 2024; Sun et al., 2024) adopted similar strategies of using large language models to augment captions. While weak-captioning does improve downstream metrics, it is suboptimal because it fails to incorporate the audio signal itself. A recent work (Huang et al., 2023b) explored a knowledge distillation approach that leverages data labels and a pre-trained audio captioner and retriever to improve caption quality. Chen et al. (2020) attempted to extract audio clips from videos by employing classifiers to detect ambient audio, speech, and music. In this work, we introduce an efficient dataset collection pipeline that relies on video datasets to extract ambient audio clips. We use this approach to collect $57M$ audio clips and use our state-of-the-art captioning method to add audio-aligned text descriptions, compromising the largest available text-audio-video dataset.
@@
Figure 1: (Left) Overview of AutoCap. We employ frozen CLAP (Kim et al., 2024b) and HTSAT (Chen et al., 2022) encoders to produce the audio representation. We then compact this representation into 4x fewer tokens using a Q-Former (Li et al., 2023a) module. This representation, along with tokens derived from pertinent metadata, is processed by a pretrained BART encoder-decoder model to generate the final caption. (Right) Overview of GenAu. Following latent diffusion models, we use a frozen 1D-VAE to convert a Mel-Spectrogram into latent sequences, which are then divided into groups and processed using ‘local’ attention layers based on the FIT architecture (Chen & Li, 2023). ‘Read’ and ‘write’ layers, implemented as cross-attention, facilitate information transfer between input latents and learnable latent tokens. Finally, ‘global’ attention layers on latent tokens allow for global communication across all groups.
@@ 3 Method
In this section, we describe our approach to high-quality text-to-audio generation, starting with audio captioning using AutoCap in section 3.1, data collection and processing in section 3.2, and ambient audio generation with GenAu in section 3.3
@@ 3.1 Automatic Audio Captioning
Audio is an inherently ambiguous modality, as many events can produce similar sound effects—a phenomenon often leveraged in animation, where soundscapes are artificially constructed. AAC attempts to generate textual descriptions for audio clips. Previous AAC methods have generally adopted an encoder-decoder transformer design, where an audio encoder is responsible for producing a representation that is processed by the decoder to produce a caption. Recent state-of-the-art methods (Étienne Labbé et al., 2023; Kim et al., 2024b) employ a pretrained audio encoder and finetune a pre-trained language model as the decoder, relying solely on the audio content for captioning. We believe that this approach is suboptimal. By directly finetuning the pre-trained language model on the limited available dataset, these methods often suffer from overfitting and limited expressiveness and accuracy. Audio files from many sources, however, are still commonly associated with metadata that might be relevant for captioning such as raw user descriptions, or a related modality (i.e. accompanied visual information). Motivated by this observation, we propose AutoCap, an audio captioning model that employs an intermediate audio representation to connect the pretrained encoder and decoder, and uses metadata to aid with the audio captioning. Figure 1 (left) presents an overview of AutoCap.
We consider a dataset of audio signals paired with a corresponding caption $\langle\mathbf{a},\mathbf{y}\rangle$ and metadata represented as a set of token sequences $\{\mathbf{m}_{j}\}_{j=1}^{j=M}$. Inspired by state-of-the-art AAC methods (Mei et al., 2023a; Étienne Labbé et al., 2023; Kim et al., 2024b), we employ an encoder-decoder architecture. We start by computing a global feature representation of the audio:
@@
where $\mathcal{P}_{\mathrm{clap}}$ is a learnable projection layer and $\mathcal{E}_{\mathrm{clap}}$ is the audio encoder of a pretrained CLAP model (Elizalde et al., 2023). Then we compute a local feature representation of the input audio:
@@
where $\mathcal{Q}$ is a Q-Former (Li et al., 2023a) that outputs a compact sequence audio representation and $\mathcal{E}_{\mathrm{a}}$ is a pretrained HTSAT (Chen et al., 2022) audio encoder that produces a time-aligned representation. The Q-Former efficiently learns 256 latent tokens, which serve as keys in cross-attention layers with the input features, thereby condensing the audio input features into 256 tokens. Metadata sequences $\mathbf{m}_{i}$ are then embedded using the embedding layer of the pretrained decoder model to obtain corresponding embedding sequences $\mathbf{x}_{\mathrm{meta}_{i}}$. For our experiments, we use video titles and captions as the metadata. We represent the input audio and metadata as the following input sequence:
$\mathbf{x}=\mathbf{x}_{\mathrm{clap}}~{}\texttt{[boa]}~{}\mathbf{x}_{\mathrm{audio}}~{}\texttt{[eoa]}~{}\texttt{[bom]}_{1}~{}\mathbf{x}_{\mathrm{meta}_{1}}~{}\texttt{[bom]}_{1}~{}...~{}\texttt{[bom]}_{M}~{}\mathbf{x}_{\mathrm{meta}_{M}}~{}\texttt{[bom]}_{M}\text{,}$ (3)
@@
$\mathbf{\hat{y}}=\mathcal{D}_{\mathrm{t}}(\mathbf{x})\text{,}$ (4)
where $\mathcal{D}_{\mathrm{t}}$ is a pretrained BART transformer model (Lewis et al., 2020) serving as the decoder. Finally, we train our model using a standard cross-entropy loss over next token predictions: